### v2

# Algebraic Limits of Sandpiles

Abstract
The paper contributes to building algebraic foundations of self-organized criticality answering a previously unsolved question about the limiting structure of the extended sandpile group as well as relating it to another limit at the level of classical sandpile groups with respect to certain monomorphisms, and puts forward a concept of canonical sandpile epimorphisms, drawing an unexpected consequence about the divisibility properties of the numbers of spanning trees on rectangles.
1 Introduction
What we now call “sandpile model” has appeared multiple times in diverse guises throughout mathematics and physics, but has acquired a tremendous attention of a large scientific community only when it was proposed as a prototype for self-organized criticality [1]. The name of the model used in the present note was coined by Deepak Dhar with an additional adjective “Abelian” referring in part to his discovery [2] of the fact that the set of recurrent states of the model on a finite domain has a natural structure of an Abelian group.
One characteristic feature of self-organized critical systems is their scale transcendence, which manifests in the form of various scaling limit theorems such as [3] or [4]. These type of results may be described as geometric or “visual” approach to the limiting structure of sandpiles, since the graphical representation of states is under consideration. Alternatively, one may focus on probabilistic properties of the sandpile model in the scaling limit (see for instance [5]) ignoring altogether how the states look like.
What is going to be discussed next is similar to the latter spirit, but the attention will be turned to the algebraic aspects of the theory, i.e. our goal is to prepare the ground for the abstract understanding of sandpile groups in the limit, ignoring for the moment the intricate structure and stunning beauty of the underlying recurrent configurations. Perhaps, over time, these three approaches – geometric, probabilistic and algebraic – will be unified.
2 Sandpile groups and harmonic functions
Consider a finite connected undirected graph without loops with a distinguished vertex that we call the sink. Denote by $\Gamma$ the set of non-sink vertices of this graph. A state of the sandpile model is a non-negative integer-valued function on $\Gamma.$ One may think that a value of a state at a vertex represents the number of sand grains.
The (reduced) Laplacian $\Delta\colon\mathbb{Z}^{\Gamma}\rightarrow\mathbb{Z}^{\Gamma}$ is given by the difference of the symmetric adjacency matrix and the diagonal degree matrix with removed row and column corresponding to the sink. Note that the graph is restored from its Laplacian, the absolute value of its determinant is equal to the number of spanning trees on the graph by Kirchhoff’s theorem.
A state $\phi$ is called stable if $\phi(v)$ is less than the degree of $v$ for all $v\in\Gamma.$ If a state is unstable there exist a vertex $v$ where $\phi(v)$ is greater or equal to the degree of $v$ and a toppling (i.e. redistribution of sand grains) $\phi\mapsto\phi+\Delta\delta_{v}$ can be performed. Here $\delta_{v}$ is a function on $\Gamma$ equal to $1$ at $v$ and vanishing otherwise. A relaxation is a process of applying topplings until the state becomes stable. The process is always finite and the resulting state $\phi^{\circ}$ doesn’t depend on the order of topplings.
A stable state is called recurrent if it can be obtained from any other state by adding sand grains and relaxing. The sandpile group $G(\Gamma)$ consists of recurrent states with an operation given by $(\phi,\psi)\mapsto(\phi+\psi)^{\circ}.$ For a function $f\in\mathbb{Z}^{\Gamma}$ there exist a unique recurrent state $\phi$ and a function $H\in\mathbb{Z}^{\Gamma}$ such that $f=\phi+\Delta H.$ This gives an isomorphism $G(\Gamma)\cong\operatorname{Coker}(\Delta)=\mathbb{Z}^{\Gamma}/\Delta(\mathbb{Z}^{\Gamma})$ which can be written equivalently in a form of a short exact sequence
$0\rightarrow\mathbb{Z}^{\Gamma}\xrightarrow{\Delta}\mathbb{Z}^{\Gamma}\rightarrow G(\Gamma)\rightarrow 0,$
where the notaion means that the left homomorphism is injective (which follows from the non-vanishing of the determinant of $\Delta$), the right one is surjective, and the image and kernel coincide in the middle.
Denote by $\partial\Gamma\subset\Gamma$ the set of vertices adjacent to the sink. We call $\partial\Gamma$ the boundary and $\Gamma^{\circ}=\Gamma\backslash\partial\Gamma$ the interior of the graph. The extended sandpile group $\widetilde{G}(\Gamma)$ is derived by allowing the recurrent states to take real non-negative values along $\partial\Gamma$. In other words, there is a short exact sequence
$0\rightarrow\mathbb{Z}^{\Gamma}\xrightarrow{\widetilde{\Delta}}\mathbb{Z}^{\Gamma^{\circ}}\mkern-9.0mu\oplus\mathbb{R}^{\partial\Gamma}\rightarrow\widetilde{G}(\Gamma)\rightarrow 0,$
where $\widetilde{\Delta}$ is the composition of $\Delta$ with the inclusion of $\mathbb{Z}^{\Gamma}=\mathbb{Z}^{\Gamma^{\circ}}\mkern-9.0mu\oplus\mathbb{Z}^{\partial\Gamma}$ to $\mathbb{Z}^{\Gamma^{\circ}}\mkern-9.0mu\oplus\mathbb{R}^{\partial\Gamma}.$ There is another short exact sequence
$0\rightarrow G(\Gamma)\rightarrow\widetilde{G}(\Gamma)\xrightarrow{/\mathbb{Z}}(\mathbb{R}/\mathbb{Z})^{\partial\Gamma}\rightarrow 0,$
where $/\mathbb{Z}$ takes the fractional part of the extended recurrent state and $\mathbb{R}/\mathbb{Z}$ is the quotient of $\mathbb{R}$ by the additive subgroup $\mathbb{Z}$ which topologically is a circle.
Let $\Delta^{\circ}\colon\mathbb{Z}^{\Gamma}\rightarrow\mathbb{Z}^{\Gamma^{\circ}}$ be a map given by the composition of $\Delta$ with the restriction to $\Gamma^{\circ}.$ For an Abelian group $A$ define the group $\mathcal{H}_{A}(\Gamma)$ of $A$-valued harmonic functions as the kernel of $\Delta^{\circ}\otimes A.$
Consider a category $\mathcal{D}\mkern-12.0mu\mathcal{D}$ of connected graphs without loops (possibly infinite, with or without sink) represented by pairs $(\Gamma,\Delta)$ with morphisms given by embeddings $\iota\colon\Gamma_{1}\rightarrow\Gamma_{2}$ such that $\Delta_{1}$ is obtained from $\Delta_{2}$ by removing rows and columns corresponding to vertices in $\Gamma_{2}\backslash\iota(\Gamma_{1}).$ For example, the casual map $\Omega\subset\mathbb{R}^{n}\mapsto\Omega\cap\mathbb{Z}^{n}$ is upgraded to a contravariant functor from the category of open sets in $\mathbb{R}^{n}$ with inclusions as morphisms to $\mathcal{D}\mkern-12.0mu\mathcal{D}.$ Note that $A^{\Gamma}$ and $A^{\Gamma^{\circ}}$ are contravariant functors from $\mathcal{D}\mkern-12.0mu\mathcal{D}$ to Abelian groups and $\Delta^{\circ}\otimes A$ is their natural transformation. Therefore, $\mathcal{H}_{A}(\Gamma)$ is also a contravariant functor, with the induced homomorphisms being simple restrictions.
Proposition 1.
The groups $\widetilde{G}(\Gamma)$ and $\mathcal{H}_{\mathbb{R}/\mathbb{Z}}(\Gamma)$ are canonically isomorphic.
Proof.
The isomorphism is given by the connecting homomorphism arising in the following employment of the snake lemma where for the middle row we use the invertibility of $\Delta\otimes\mathbb{R}:$
[…]
Remark 1.
In a similar way, one obtains $G(\Gamma)\cong\operatorname{Ker}(\Delta\otimes\mathbb{R}/\mathbb{Z}),$ the latter may be thought of as the group of “strictly harmonic” circle-valued functions on the graph, i.e. the vanishing of the Laplacian is required both in the interior and on the boundary.
Another application of the snake lemma gives the following.
Proposition 2.
The groups of integer-, real- and circle- valued harmonic functions and the group $\operatorname{Coker}\Delta^{\circ}$ are linked through an exact sequence
$0\rightarrow\mathcal{H}_{\mathbb{Z}}(\Gamma)\rightarrow\mathcal{H}_{\mathbb{R}}(\Gamma)\rightarrow\mathcal{H}_{\mathbb{R}/\mathbb{Z}}(\Gamma)\rightarrow\operatorname{Coker}\Delta^{\circ}\rightarrow 0$
Lie groups $\mathcal{H}_{\mathbb{R}}(\Gamma)$ and $\mathcal{H}_{\mathbb{R}/\mathbb{Z}}(\Gamma)$ both have dimension equal to $|\partial\Gamma|$ and $\operatorname{Coker}\Delta^{\circ}$ is discrete. Therefore, since $\mathcal{H}_{\mathbb{R}}(\Gamma)$ is connected (being an $\mathbb{R}$-vector space), the image of $\mathcal{H}_{\mathbb{R}}(\Gamma)$ in $\mathcal{H}_{\mathbb{R}/\mathbb{Z}}(\Gamma)$ is identified with the maximal connected subgroup of $\widetilde{G}(\Gamma)$ that we denote by $\widetilde{G}_{0}(\Gamma).$ We observe the isomorphisms $\widetilde{G}_{0}(\Gamma)\cong\mathcal{H}_{\mathbb{R}}(\Gamma)/\mathcal{H}_{\mathbb{Z}}(\Gamma)$ and $\operatorname{Coker}\Delta^{\circ}\cong\widetilde{G}(\Gamma)/\widetilde{G}_{0}(\Gamma).$
Yet another snake lemma diagram (which is also omitted) explains the relation between the cokernel of $\Delta^{\circ}$ and the usual sandpile group.
Proposition 3.
There is an exact sequence
$0\rightarrow\mathcal{H}_{\mathbb{Z}}(\Gamma)\rightarrow\mathbb{Z}^{\partial\Gamma}\rightarrow G(\Gamma)\rightarrow\operatorname{Coker}\Delta^{\circ}\rightarrow 0$
Here, the map $\mathcal{H}_{\mathbb{Z}}(\Gamma)\rightarrow\mathbb{Z}^{\partial\Gamma}$ is the composition of $\Delta$ and the projection to $\partial\Gamma.$ Denote by $G_{0}(\Gamma)$ the image of the homomorphism $\mathbb{Z}^{\partial\Gamma}\rightarrow G(\Gamma).$ It is a subroup of the sandpile group generated by adding sand at the boundary. Note that $G_{0}(\Gamma)=\widetilde{G}_{0}(\Gamma)\cap G(\Gamma)$ and $\operatorname{Coker}\Delta^{\circ}\cong G(\Gamma)/G_{0}(\Gamma).$ This implies the following lemma.
Lemma 1.
If $G(\Gamma)$ is generated by adding sand at the boundary, then
$\widetilde{G}(\Gamma)\cong\mathcal{H}_{\mathbb{R}}(\Gamma)/\mathcal{H}_{\mathbb{Z}}(\Gamma).$
For the rest of this section we will be concerned with the inverse limit
$\lim_{\mathbb{Z}^{2}\leftarrow\Gamma}\widetilde{G}(\Gamma)$
over all finite $\Gamma\subset\mathbb{Z}^{2}.$ It can be replaced with a limit over any sequence of finite subgraphs $\Gamma_{1}\subset\Gamma_{2}\subset\dots\subset\mathbb{Z}^{2}$ such that $\cup_{n=1}^{\infty}\Gamma_{n}=\mathbb{Z}^{2}.$ We prefer to assume that all $\Gamma_{n}$ are convex (i.e. intersections of convex domains with the lattice) for the following reason.
Lemma 2.
For convex $\Gamma_{1}\subset\Gamma_{2}\subset\mathbb{Z}^{2}$ the restriction homomorphism
$\mathcal{H}_{A}(\Gamma_{2})\rightarrow\mathcal{H}_{A}(\Gamma_{1})$
is surjective.
Proof.
One can approach this by observing first that the restriction homomorphisms to the domain from its diamond hull (the minimal diamond, not necessarily Aztec, containing this domain) is an isomorphism. The surjectivity for two diamonds, where one is inside the other, is obvious. ∎
Lemma 3.
For $\Gamma\subset\mathbb{Z}^{2}$ the sandpile group $G(\Gamma)$ is generated by adding sand at the boundary.
Proof.
By induction one can show that adding a grain of sand at a vertex can be achieved by adding grains to the left of it and at the boundary. ∎
Lemma 4.
For a sequence of Abelian groups and epimorphisms
$A_{1}\xleftarrow{\pi_{1}}A_{2}\xleftarrow{\pi_{2}}A_{3}\xleftarrow{\pi_{3}}\cdots$
there is a short exact sequence
$0\rightarrow\lim_{\leftarrow}A_{n}\rightarrow\prod_{n=1}^{\infty}A_{n}\xrightarrow{\Psi}\prod_{n=1}^{\infty}A_{n}\rightarrow 0,$
where $\lim_{\leftarrow}A_{n}$ is the inverse limit, $\prod_{n=1}^{\infty}A_{n}$ is the direct product and
$\Psi(a_{1},a_{2},\dots)=(\pi_{1}(a_{2})-a_{1},\pi_{2}(a_{3})-a_{2},\dots).$
Proof.
The inverse limit is by definition the kernel of $\Psi.$ The surjectivity is implied by surjectivity of $\pi_{1},\pi_{2}\dots$
Theorem 1 (Affirmative answer to Question 1 from [6]).
The natural inclusion of $\mathcal{H}_{\mathbb{R}}(\mathbb{Z}^{2})/\mathcal{H}_{\mathbb{Z}}(\mathbb{Z}^{2})$ to $\lim_{\mathbb{Z}^{2}\leftarrow\Gamma}\widetilde{G}(\Gamma)$ is an isomorphism.
Proof.
Note that for any Abelian group $A$
$\mathcal{H}_{A}(\mathbb{Z}^{2})\cong\lim_{\mathbb{Z}^{2}\leftarrow\Gamma}\mathcal{H}_{A}(\Gamma).$
Lemmas 3 and 1 imply exactness of columns, and Lemmas 2 and 4 give exactness of rows in the diagram below, now we can apply (for the fifths and last time) the snake lemma:
00{0}00{0}00{0}${\mathcal{H}_{\mathbb{Z}}(\mathbb{Z}^{2})}$${\prod\mathcal{H}_{\mathbb{Z}}(\Gamma_{n})}$${\prod\mathcal{H}_{\mathbb{Z}}(\Gamma_{n})}$00{0}${\mathcal{H}_{\mathbb{R}}(\mathbb{Z}^{2})}$${\prod\mathcal{H}_{\mathbb{R}}(\Gamma_{n})}$${\prod\mathcal{H}_{\mathbb{R}}(\Gamma_{n})}$00{0}${\mathcal{H}_{\mathbb{R}/\mathbb{Z}}(\mathbb{Z}^{2})}$${\prod\mathcal{H}_{\mathbb{R}/\mathbb{Z}}(\Gamma_{n})}$${\prod\mathcal{H}_{\mathbb{R}/\mathbb{Z}}(\Gamma_{n})}$00{0}00{0}00{0}
3 Monomorphisms and epimorphisms
For an Abelian group $A$ denote by $A^{*}$ its Pontryagin dual, i.e.
$A^{*}=\operatorname{Hom}(A,\mathbb{R}/\mathbb{Z}).$
In general, if $A$ is finite then it is isomorphic to $A^{*}.$ However, a choice of such an isomorphism is equivalent to an extra structure on $A$ which is a non-degenerate paring on $A$ with values in the circle $\mathbb{R}/\mathbb{Z}.$ For a sandpile group of a graph such a paring comes naturally.
Proposition 4.
Sandpile groups of finite graphs are canonically self-dual.
Proof.
To construct the pairing on $G(\Gamma)$ we apply natural isomorphisms $G(\Gamma)\cong\operatorname{Coker}\Delta$ and $G(\Gamma)\cong\operatorname{Ker}(\Delta\otimes\mathbb{R}/\mathbb{Z})$ to the components of the pairing
$Q\colon\operatorname{Coker}\Delta\times\operatorname{Ker}(\Delta\otimes\mathbb{R}/\mathbb{Z})\rightarrow\mathbb{R}/\mathbb{Z}$
given by
$Q(\phi+\Delta\mathbb{Z}^{\Gamma},\psi)=\sum_{v\in\Gamma}\phi(v)\psi(v).$
First, we need to show that $Q$ is well defined, i.e. the right hand side doesn’t depend on the choice of a representative $\phi$ in the coset. This follows from observing
$\sum_{v\in\Gamma}(\Delta\delta_{w})(v)\psi(v)=(\Delta\psi)(w)=0$
for any $w\in\Gamma.$ Second, we need to show that the homomorphism
$\psi\in\operatorname{Ker}(\Delta\otimes\mathbb{R}/\mathbb{Z})\mapsto Q(-,\psi)\in(\operatorname{Coker}\Delta)^{*}$
is injective (and therefore bijective due to the same cardinality). Indeed, its kernel is trivial since for non-zero $\psi$ there exist $w\in\Gamma$ such that $\psi(w)\neq 0$ and so $Q(\delta_{w}+\Delta\mathbb{Z}^{\Gamma},\psi)=\psi(w)\neq 0.$
This proposition (and the vanishing of $\operatorname{Ext}^{1}(A,\mathbb{R}/\mathbb{Z})$ for a finite Abelian group $A$) implies that there is a canonical one-to-one correspondence between monomorphisms from $G(\Gamma_{1})$ to $G(\Gamma_{2})$ and epimorphisms from $G(\Gamma_{2})$ to $G(\Gamma_{1})$ for any two finite graphs.
A particular set of monomorphisms between some sandpile groups was introduced in [7] in the case of certain types of tilings. The construction presented there is quite involved, thus it seems to be appropriate to write below a more direct definition of these monomorphisms in the simplest case of a rectangular tiling.
Let $\Gamma(p,q)$ be the portion of the square lattice given by the intersection with an open rectangle $(0,p)\times(0,q),$ where $p$ and $q$ are positive integers. Think of elements in $G(\Gamma(p,q))$ as of strictly harmonic circle valued functions on $\Gamma(p,q)$ as in Remark 1. Then, for any pair $(m,n)$ of positive integers, the monomorphism from $G(\Gamma(p,q))$ to $G(\Gamma(mp,nq))$ is given by assigning to $\phi$
$(x,y)\in\Gamma(mp,nq)\mapsto(-1)^{\lfloor\frac{x}{p}\rfloor+\lfloor\frac{y}{q}\rfloor}\phi(\mu^{-1}_{p,\lfloor\frac{x}{p}\rfloor}(x),\mu^{-1}_{q,\lfloor\frac{y}{q}\rfloor}(y)),$
[…]
$\mu_{p,k}(x)=\begin{cases}kp+x,&\text{if $k$ even}\\ (k+1)p-x,&\text{otherwise.}\end{cases}$
The dual map (which by above mentioned considerations is an epimorphism) is given by assigning to an element $\psi$ of $G(\Gamma(mp,nq))$ a circle-valued function
$(x,y)\in\Gamma(p,q)\mapsto\sum_{k=0}^{m-1}\sum_{l=0}^{n-1}(-1)^{k+l}\psi(\mu_{p,k}(x),\mu_{q,l}(y)).$
Next we observe that the composition
$G(\Gamma(p,q))\hookrightarrow G(\Gamma(mp,nq))\twoheadrightarrow G(\Gamma(p,q))$
of the epimorphism and the monomorphism is the identity homomorphism multiplied by $mn$. There are two extreme cases: when $mn$ is coprime with $\gamma(p,q)=|G(\Gamma(p,q))|,$ in which case $G(\Gamma(p,q))$ is realized as a direct summand of $G(\Gamma(mp,nq)),$ or when $mn$ is divisible by $\gamma(m,n),$ and therefore we realize the group $G(\Gamma(p,q))$ as a subgroup of the kernel of the epimorphism, which gives the second assertion of the following theorem.
$m$ $n$ 2 3 4 5 6
2 4 15 56 209 780
3 15 192 2415 30305 380160
4 56 2415 100352 4140081 170537640
5 209 30305 4140081 557568000 74795194705
6 780 380160 170537640 74795194705 32565539635200
Table 1: Orders of sandpile groups on $(0,m)\times(0,n)\cap\mathbb{Z}^{2}.$ Observe the divisibility properties in accordance with Theorem 2. For instance, $\gamma(2,4)$ is divisible by $\gamma(2,2),$ but not by $(\gamma(2,2))^{2}$ since the scaling factor $2=1\times 2$ is not 00 modulo $\gamma(2,2)=4;$ on the other hand $(\gamma(2,2))^{2}$ divides $\gamma(4,4).$
Theorem 2.
Let $\gamma(p,q)=|G(\Gamma(p,q))|$ be the number of spanning trees on the rectangular graph with non-sink vertices $\Gamma(p,q).$ Then for any quadruple of positive integers $n,m,p,q$ the number $\gamma(p,q)$ divides the number $\gamma(mp,nq).$ Moreover, if $mn$ is divisible by $\gamma(p,q)$, then $(\gamma(p,q))^{2}$ divides $\gamma(mp,nq).$
Now we would like to describe the relationship between the sandpile monomorphisms and the restriction homomorphisms between the corresponding extended sandpile groups. Namely, the monomorphism may be seen as a section of the restriction defined over the usual sandpile group.
Proposition 5.
There is the following commutative diagram:
${G(\Gamma(m,n))}$${G(\Gamma(pm,qn))}$${\widetilde{G}(\Gamma(m,n))}$${\widetilde{G}(\Gamma(pm,qn)).}$
A particular corollary of this proposition (the proof of which is trivial) is that the direct limit of sandpile groups of rectangular domains with respect to our monomorphisms embeds to the inverse limit of extended sandpile groups which was studied in the previous section. The main problem we put forward is to identify the image of this embedding. For instance, one may expect that it is a rational subtorus or a divisible group (a step in this direction was previously made in [7], where it was shown that the limit contains elements of all finite orders). This expectation translates nicely in the dual language, i.e. that the inverse limit of (usual) sandpile groups of rectangular domains under the epimorphisms has no torsion.
It is instructive to see what is the situtation in the one-dimensional case, which is as a general rule dramatically simpler. The extended sandpile group of an integer interval $(0,n)\cap\mathbb{Z}$ for $n\geq 3$ is a two-dimensional torus made of circle-valued arithmetic progressions, and the usual sandpile group is a cyclic group of order $n,$ corresponding to such progressions which take $0\in\mathbb{R}/\mathbb{Z}$ at 00 and $n.$ The direct limit of this groups is $\mathbb{Q}/\mathbb{Z},$ a rational circle corresponding to periodic circle-valued arithmetic progressions with value 00 at the origin, and the inverse limit under the epimorphisms is ^$\widehat{\mathbb{Z}}=\prod_{l}\mathbb{Z}_{l},$ a direct product of all additive groups of $l$-adic integers, which indeed has no torsion.
4 Discussion
There is a striking and not fully explored relation [8] between the sandpile model on a rectangular domain and domino tilings on it. The group playing the role in this relation is not the whole sandpile group of the rectangle, but rather a subgroup of elements symmetric with respect to horizontal and vertical reflections. It seems that with a bit of extra work our Theorem 2 may be adapted to the case of such subgroups, and therefore give the analogous divisibility result for the numbers of domino tilings.
The category $\mathcal{D}\mkern-12.0mu\mathcal{D}$ which was passingly introduced in Section 2 imitates by its essence a category of open subsets of some topological space. It makes sense to extend its class of morphisms, leaving the objects intact, by allowing all maps between graphs whose pullbacks on integer-valued functions commute with $\Delta^{\circ}.$ Morally, this allows for coverings, thus we may denote such a category as $\mathcal{D}\mkern-12.0mu\mathcal{D}^{et}$ with $et$ standing for “étale”, and $\mathcal{H}_{A}$ is still some kind of a sheaf. It is conceivable that it would be fruitful to study “the universal extended sandpile group” defined as the inverse limit (i.e. global sections) of $\mathcal{H}_{\mathbb{R}/\mathbb{Z}}$ over this category. In addition, it looks like there are traces of higher homology theory, as the preparations for the proof of Theorem 1 suggest.
Studying more general algebraic structures such as sandpile monoid (where non-recurrent states are taken into account) and sandpile magma (where negative values are allowed), as well as their real boundary extensions, is a very promising direction for further exploration. At the moment we don’t know if similar functoriality properties to that of an extended sandpile group hold, but some of the most classical entities, such as the first “sandpile movie” (which has in part inspired the introduction [6] of other movies driven by harmonic potentials) suggested by Michael Creutz in [9] connecting the zero element in the monoid to the neutral element in the group, are naturally dwelling there as geodesic paths. Moreover, it seems to be plausible that the extended sandpile magma is realized geometrically as a tropical modification of the extended sandpile group.
In conclusion, here are a few words about the relevance of tropical view. First of all, another motivation for the idea of extending the sandpile model has arisen from the following interpretation of the scaling limit theorems stated (in their weaker form) in [4], and with their consequences drawn in [10]: infinitesimal behavior of the sandpile group in high-energy regime near the scaling limit is described by tropical geometry. Since the usual sandpile group is discrete, this sentence is hard to comprehend literally. However, the extension at the boundary, brings the continuous structure, and the extended sandpile group is indeed a tropical Abelian variety. More, this variety may be said to be defined over $\mathbb{Z},$ and its set of integral points is precisely the classical sandpile group. Nevertheless, the corresponding restriction homomorphisms are not defined over integers, and thus there is no functionality under domain inclusions at the level of classical sandpile groups – this is why the presence of our mono/epi-morphisms in some special situations is so exciting.